Steward INC latest eps was $3.50 and it’s book value per share was $22.75, it had 215,000 shares outstanding, and its debt-to-asset ratio was 46%. How much debt was outstanding ?
Answer #1

Book Value = Book Value Per Shares * Number of Shares = 22.75*215,000 = 4,891,250

DEBT to Asset Ratio = DEBT/(Debt + Equity ) = 46%
Debt/(Debt + 4891250) = 46%
Debt = 46%*Debt + 4891250*46%
Debt outstanding= 4,166,620.37

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well